# Visualization of Electronic Health Record Sequences at Scale

### v1

@@
Abstract
We present ParcoursVis, a Progressive Visual Analytics tool designed to explore electronic health record sequences of patients at scale. Existing tools process and aggregate the whole dataset upfront before showing the visualization, taking a time proportional to the data size. Therefore, to remain interactive, existing tools are limited to data sizes that can be processed in under a few seconds to meet the latency constraints of human attention. To overcome this limitation and scale to larger sizes, ParcoursVis relies on a progressive algorithm that quickly shows an approximate initial result of the aggregation, visualized as an Icicle tree, and improves it iteratively, updating the visualization until the whole computation is done. With its architecture, ParcoursVis remains interactive while visualizing the sequences of tens of millions of patients, each described with thousands of events; three to five orders of magnitude more than similar systems. Managing large datasets allows for exploring rare medical conditions or unexpected patient pathways, contributing to improving treatments. We describe the algorithms we use and our evaluation concerning their scalability, convergence, and stability. We also report on a set of guidelines to support visualization designers in developing scalable progressive systems. ParcoursVis already allows practitioners to perform analyses on two large real medical datasets. Our prototype is open-source.
keywords:
@@ Abstract
[…]
@@
The ParcoursVis system, designed to explore electronic health record sequences of patients at scale. Visualization of Electronic Health Record Sequences at Scale the main view, shows 10 million medical pathways for patients treated for non-cancerous prostate adenoma, visualized as an Icicle tree. Visualization of Electronic Health Record Sequences at Scale shows the number of filtered out and total patients. Visualization of Electronic Health Record Sequences at Scale shows, in the control panel, the legend and overall distribution of event types. Visualization of Electronic Health Record Sequences at Scale shows the hierarchy of event types used to summarize the tree. Visualization of Electronic Health Record Sequences at Scale zooms into a sub-tree, revealing more of the hierarchical structure. Each colored node represents a stage in a sequence of events starting from the root on the left, with time flowing to the right. For each node, the next event nodes are on the right, sorted from bottom to top by frequency. The node height is proportional to the frequency of the event, and the width to the average duration. In this dataset, most patients started with the “alphabloc” treatment (bottom left), and kept taking it for life (a few patients changed treatment since the node’s right side is almost empty). {subcaptiongroup} $phantomcaption \phantomcaption \phantomcaption \phantomcaption$
1 Introduction
As datasets continue to increase in size and complexity, traditional visualization techniques struggle to maintain interactivity and responsiveness, thereby hindering effective exploratory data analysis. Addressing the scalability challenge necessitates progress in algorithmic implementations and the development of optimized data structures within visualization systems, but more importantly, a change in programming paradigm, using progressive visual analytics (PVA) to decouple data size and algorithm complexity from latency [13].
The scalability challenge is particularly acute in healthcare, where Electronic Health Records (EHRs) contain vast amounts of patient information. For instance, the French National Health Database [41] holds decades of health data on citizens, encompassing records of drug purchases, medical diagnoses, and hospital treatments. Such data have the potential to advance healthcare by enabling a better understanding of treatment practices and patient outcomes. As an example, by reconstructing and visualizing temporal event sequences from drug purchase records, the examination of patients’ treatment journeys over time is facilitated. This approach aligns with the goal of medical professionals to assess whether patients’ treatments adhere to national guidelines and, if deviations occur, to understand the reasons. This is key to providing patients with the highest quality of care through evidence-based and safe recommendations. Yet, the volume of such datasets necessitates innovative visualization approaches capable of handling large-scale medical data while preserving analytical efficiency and usability. In this context, we developed ParcoursVis (Visualization of Electronic Health Record Sequences at Scale), a PVA tool created in collaboration with medical institutions to visualize patient care pathways derived from EHRs.
Existing work has laid a foundation for similar visualization tools aimed at healthcare analysis (e.g., [5, 28, 31]). For instance, EventFlow [29] aggregates event sequences into a prefix tree and visualizes them as an Icicle tree, as shown in Visualization of Electronic Health Record Sequences at Scale; a visualization that has proven effective in identifying optimal drug usage patterns. We, therefore, build on the EventFlow visualization technique for ParcoursVis, along with several improvements.
However, tools like EventFlow have generally been applied to relatively small datasets, sometimes encompassing fewer than 100 patients. According to our experiments, EventFlow is limited to about 20k patients. While such tools have offered valuable initial insights, they fall short in scalability, which has been reported as a main challenge for EHR visualization [49]. In practice, regional and national EHR databases present data at a far greater scale, often involving millions of patients. To our knowledge, no existing tool enables the management and interactive analysis of EHR datasets exceeding 100k event sequences [51, 50, 29, 49], which limits the scope of exploration. This constraint impedes the ability to gain a more comprehensive view of patient pathways, identify patterns and trends that may not be visible in smaller or sampled datasets, and discover rare patterns and outliers. With ParcoursVis, we allow visualizing temporal event sequences of more than 100M patients, potentially supporting the largest EHRs.
Large-scale interactive data processing poses significant challenges, particularly in terms of response times. To accommodate the need for a responsive system that supports health analysts in conducting explorations on large-scale EHR data, we rely on a Progressive Visual Analytics architecture [42, 13]. Our progressive approach provides a quick, approximate, yet useful preview of the data early on to maintain user attention within human latency limits and quick updates converging within seconds. Indeed, results from research in visualization show that user attention declines after 500 ms, and after five to ten seconds, users tend to abandon tasks or lose focus [52, 25]. After the initial preview, our algorithm continues processing the dataset, updating the visualization every second or so until it is fully processed.
Yet, the progressive updates of the aggregated tree can cause visual instability, as the nodes (i.e., rectangles on Visualization of Electronic Health Record Sequences at Scale) are laid out at each update. The shifting of nodes can disrupt the viewer’s mental map and create distracting visual flickers. Therefore, our approach may have an impact on the usability of the tool. To mitigate this issue, we introduce a sorting with hysteresis algorithm that limits this flicker during the progressive rendering.
In summary, this article presents our system and reports on its usability in terms of PVA, stability, and possibly flicker. Precisely, we:
  1. 1.
    Introduce our algorithm designed to process events progressively and evaluate its efficiency using metrics specific to PVA [38]. In particular, we report on scalability showing that our aggregation algorithm has linear complexity relative to the number of patients, that the processing of events per patient scales linearly with the number of CPU cores, and that it consistently meets the specified latency constraints [52, 25]. We also report on accuracy and show that ParcoursVis’s progressive aggregation of sequences converges quickly for nodes close to the root of the tree, in line with analysts who explore the tree from the root to the leaves.
  2. 2.
    Introduce a sorting algorithm with hysteresis, i.e., maintaining node order for size differences below a threshold. We report on stability, demonstrating that sorting with hysteresis enhances both stability (i.e., avoiding flickering) and convergence of the progressive visualization.
  3. 3.
    We report on usability regarding the progressive aspect. Our professional participants have completed a variety of tasks seamlessly, without mentioning any delays in system’s response or uncomfort with the progressive rendering of the Icicle tree.
According to the survey of Ulmer et al. [44], we describe the first progressive tree visualization algorithm. We outline its specific issues related to stability and uncertainty.
ParcoursVis improves the scalability of aggregated EHR event sequence visualization by at least three orders of magnitude compared to state-of-the-art systems. Its source code and evaluation scripts are available at https://gitlab.inria.fr/aviz/parcoursvis/.
2 Related Work
Our work relates to temporal event sequences, scalability, PVA systems, and user uncertainty in the partial results. Our research goal is not to propose a new visualization technique but rather to explore the strategies to scale and adapt existing approaches that have already demonstrated their effectiveness and apply them to the largest EHR data available.
2.1 Temporal Event Sequences
Our tool, ParcoursVis, aims to support domain experts in visualizing and exploring patients’ care pathways. Like EventFlow [29], it uses aggregation and filtering to visualize event sequences. According to the taxonomy in Wang et al.’s survey of EHR visualization [49], it falls into the “Event Sequence Simplification” (ESS) family of techniques (see Figure 1). This family is defined as “any technique used for reducing the visual complexity of event sequences in aggregated display overviews” [49]. Wang et al.distinguish seven families: ML, NLP, ESS, Geospatial Visualization, Clustering, Comparison, and Others. Among the 51 articles related to the visual analysis of EHR data, 11 belong to the ESS family. In other families, visualizing an overview of the event sequences is not the focus; they rather focus on one patient, a set of patients, or spatial areas sharing some characteristics.
$a_{2}$$b_{2}$$c_{1}$$d_{1}$$e_{1}$
$a_{2}$$b_{2}$$c_{1}$$d_{1}$$e_{1}$
Figure 1: ESS Aggregation of EHR Sequences as implemented by EventFlow and ParcoursVis. Patient A took the treatment sequence $\{a,b,c\}$ in that order, and patient B took $\{a,b,d,e\}$. The left side shows the prefix tree of these sequences, and the right side shows the Icicle tree visualization where node height encodes the frequency.
Visualization of Electronic Health Record Sequences at Scale shows the aggregated tree built from 10M sequences (patients) of events (e.g., treatments), visualized as an Icicle tree. As explained in Figure 1, the node height is proportional to the frequency. The seven nodes on the left of the visualization (Visualization of Electronic Health Record Sequences at Scale) represent these seven options at the beginning of the treatment. A patient can start by taking one or a mix of three drugs or directly undergo surgery. Each node has a width proportional to the duration of the treatment. The duration of treatment is often computed by aggregating several consecutive low-level events describing a patient repeatedly buying a prescribed drug at specific times, producing one continuous treatment event with a duration. ParcoursVis reuses the visualization technique of EventFlow because it was thoroughly tested and validated for analyzing medical temporal event sequences [5, 29, 31]. EventFlow, however, handles datasets with a limited number of sequences [5, 29, 31, 12]; about 20k sequences according to our experiments (available in the Supplemental Material). A common strategy for handling larger datasets is to select a random sampling of sequences [12]. Random sampling, however, loses rare and, therefore, important sequences that interest practitioners. Instead, we use PVA to be scalable and accurate.
In addition to visualization, EHR data is also analyzed using sequence mining augmented with visualization [46, 27, 33, 42]. The idea consists of trying to automatically discover frequent and important patterns in the sequences to support the analyst in discovering unexpected information. This approach differs from ours since it only needs to visualize the search results. Instead, our tool interactively visualizes all the treatment pathways following Shneiderman’s mantra: overview first, zoom, filter, and details on demand [40].
2.2 Scalability
Existing systems that analyze event sequences, both in medical and non-medical domains, often fail to report their scalability [38] relative to the dataset sizes they handle. Wang et al.’s [49] survey does not report dataset sizes, although it reports scalability as the first future research challenge. In the medical domain, reported dataset sizes vary widely, with studies analyzing as few as 65 patients [3] to as many as 833,710 public health cases spanning over 10 years [21], passing by EventFlow [29] and its few thousand patients. Some works, such as DecisionFlow [18], mention handling over a million individual point events but provide no details on computational efficiency. In contrast, many other studies work with significantly smaller datasets, such as 5,800 patients [20], 2,336 patients [34], and 1,186 subjects [22]. The situation is similar in non-medical domains, where studies report dataset sizes in the range of thousands to millions of sequences, the largest dealing with 6,477 sequences of daily activities and 989,925 website visit sequences [46], yet offer no information on computational performance. We easily visualize this dataset with ParcourVis (available in Supplementary Material). Additionally, Liu et al. [26] mention that they use real-world datasets that can contain hundreds of thousands of unique sequences, but do not provide algorithmic details for performance data.
Overall, for systems dealing with event sequence data, there is either a lack of information on performance relative to dataset size or a focus on small datasets, making it difficult to assess scalability.
2.3 Progressive Visual Analytics
PVA allows users to visualize the result of an algorithm while it is computed, in a progressive way. Instead of processing the whole data at once, waiting an unpredictable time, as standard algorithms do, PVA algorithms process data chunk by chunk or iteratively, or a mix of both [13] to control the latency and show intermediate results iteratively at a controlled pace. The main goal is to keep the overall system interactive and yield results within the typical user’s attention span that lasts 1–10 seconds [30]. Practitioners already applied and studied the PVA paradigm on multiple types of visualizations, as described in the survey of Ulmer et al. [44].
In this work, we progressively aggregate event sequences of patients’ care pathways into a prefix tree [49] that we visualize as an Icicle tree. Although PVA has been used for pattern mining in EHRs [42], to the best of our knowledge, we are the first to apply PVA to visualize EHR aggregated data at scale. We also introduce the first progressive tree visualization technique, according to the survey of Ulmer et al.
By presenting partial results quickly, PVA systems allow users to get quick feedback on the overview and queries. It also increases users’ commitment and actions per minute [52]. In this article, we show that our PVA algorithm for computing the prefix tree is indeed scalable, according to the definition of Richer et al. [38]: the aggregation time scales linearly with the number of events, inversely proportional to the number of threads, and updates always meet the latency deadline.
2.4 Uncertainty and Trust
PVA systems start by showing partial results, implying uncertainty due to the progressive aspect of the computation. This progressive uncertainty [45] is reduced over time until the final result is computed, without any progressive uncertainty but potentially with other types of uncertainty, just like standard data and visualizations. The visualization shown early may or may not be close to the final result. It is, therefore, important for analysts to assess the quality of the progressive results to decide if they can trust the visualization or if they should wait longer.
Angelini et al. [2] explain that progressive visualizations undergo three stages: early partial results that are usually noisy and not trustworthy, mature partial results that are stabilizing and can be trusted but with uncertainty, and definitive partial results that are accurate and can take (wasted) time to finish. When our aggregation tree is computed progressively, all the nodes will undergo these three stages. Yet, the nodes closest to the root will become reliable and stable earlier than the deeper nodes since they are aggregating more sequences, and their stability is essentially related to the number of sequences they represent, assuming a well-behaved distribution of patients in nodes.
The uncertainty can be computed and visualized explicitly, through e.g.,, error bars, or implicitly through a proxy measure such as stability. If the visualization is not stable at some stage of the progression, it means that the result remains uncertain. The inverse is not always true: a progressive visualization that has remained stable for a certain time might still change before the end [32, 9].
Representing the uncertainty explicitly is possible but challenging. Understanding uncertainty in static visualization is not easy, even for statisticians [9]. For progressive visualization, many representations are possible, e.g., bar charts with error bars, violin plots, or gradient distribution [9, 15], but few have been studied and validated empirically. Patil et al. [32] have tried four designs for visualizing uncertainty on progressive bar charts, showing that two are effective. However, there is no study about the progressive visualization of an Icicle tree, to the best of our knowledge.
ParcoursVis takes a few seconds to update its tree fully, with an update every few seconds. It is not clear if users could make sense of any uncertainty visualization for many nodes in such a short update time. In our case, it might take as much time or more to make sense of an uncertainty visualization than to wait for the progressive visualization to complete. Instead, we rely on stability as a proxy for quality, assuming that subtrees will remain stable when the progressive uncertainty decreases. On the other side, when a subtree is not stabilized for a few seconds, the user will not be able to infer much from watching it. We study the convergence and stability of ParcoursVis’s progressive Icicle tree visualization; instability becomes a proxy for uncertainty [45].
3 ParcoursVis
ParcoursVis is a web-based application inspired by EventFlow that processes data in a back end; users interact via a web browser front end. While ParcoursVis is specialized in exploring and analyzing EHR sequences, its core structure is more general and can handle the same kinds of event sequences and applications as EventFlow. Contrary to EventFlow, the rules for aggregating and merging events are written in C++ for performance and expressive power reasons. For real pathologies, many constants and aggregation rules need to be set according to state-of-the-art recommendations that cannot be expressed within EventFlow. Therefore, an engineer is needed to configure ParcoursVis to define new domain-specific rules when it is applied to a new medical question with specific kinds of events.
3.1 Example Use Case and Tasks
ParcoursVis is currently used in two main applications: non-cancerous prostate adenoma treatment analysis for the French social security and analysis of emergency patient pathways in all the public Parisian hospitals; we use examples from the former. Specific to our non-cancerous prostate adenoma use case, our application receives information from the Social Security reimbursement database (SNDS) about the drugs bought by patients (time, name, molecule), dedicated surgery, and deaths. There are three molecules possibly administered for treatment: alpha-blockers, specific herbal medicine, and 5-alpha, respectively labeled alphabloc, phyto, and fivealpha in Visualization of Electronic Health Record Sequences at Scale. Those are the three concrete events our referring doctors selected to extract for their analysis. We call them the low-level events. ParcoursVis aggregates this information and synthesizes new events that we call high-level events. When the patient buys a box of pills with the same molecule every two weeks or so, all the successive low-level events are aggregated as one high-level event with a computed duration, e.g., “alphabloc” for 300 days. When the patient stops buying a box for more than six months, an “interruption” event is synthesized, and if the interruption is longer than one year, a “no treatment” event is synthesized, both with appropriate durations. Those thresholds are specified by the referring doctors according to state-of-the-art recommendations. Some of the low-level events can be combined, leading to specific aggregation rules: buying a box of alphabloc and a box of phyto within a certain time creates an “alphabloc phyto” treatment (line 3 of Visualization of Electronic Health Record Sequences at Scale). This rule only applies to the three molecules, not to other events (surgery, death); it is domain-specific. Our dataset also contains the age and information related to known chronic diseases of the patients (e.g., diabetes, hypertension) to explore possible correlations with treatment pathways.
ParcoursVis supports the following abstract tasks:
Simple Tasks that users may complete without assistance:
  • Overview: visualize all the temporal event sequences (high-level event types).
  • Zoom and Filter Events: visualize all the temporal sequences having a given prefix (children of a selected node) or filter out specific event types.
  • Details on demand: visualize all the attributes and distributions related to a specific sequence (e.g., age distribution and duration distribution of the patients reaching a node).
Advanced Tasks that may require initial guidance:
  • Filter by attribute: simplify the visualization by filtering the data based on the attributes and metadata of the sequence (e.g., its total length, age of the person it represents), and the resulting tree (e.g., size of the nodes).
  • Configure view: setting an alignment prefix to visualize what had happened before and after a specific sequence.
  • Abstracting: regrouping multiple types of events into a super-type, using a hierarchy (see Visualization of Electronic Health Record Sequences at Scale).
  • History and comparison: navigate through the history, i.e., previous filters, zooms, or alignments. Compare different trees corresponding to different history states by switching between them.
Analytical Tasks professionals want to answer:
  • Application questions: identify trends, frequent, critical, or surprising pathways, and understand underlying causes, finding actionable strategies for improvement or resolution.
  • Data questions: identify data errors, limitations, and determine more suitable data extraction methods.
3.2 User Interface
Figure 2: The “Sequence” view visualization (squared) shows all events leading to a specified prefix, here the “surgery” event. About 60% of the patients who underwent surgery were treated before. A button allows showing all events that came after the prefix instead.
Our visualizations and user interfaces (GUI) are close to EventFlow, which has already been validated through many applications and case studies [5, 29, 31]. We focus on the specific features introduced by PVA.
3.2.1 Main View and Prefix View
The “Main” view (Visualization of Electronic Health Record Sequences at Scale) shows the aggregated tree of all patients’ care pathways. The width of a node represents the mean duration of the high-level event it is associated with; its height represents the number of sequences (frequency). We sort the children of each node according to their frequency with some caveats described in Section 3.4.4 to improve the stability of successive progressive updates. The color of a node describes its high-level event type. Clicking on a node shows details on demand in the “Detail” tab of the control panel (3(a)).
Users can change the view and focus on a specific event or sequence of events. The “Prefix” view (Figure 2) shows the sequence tree that leads to or follows a given prefix, specified interactively.
Each of these views is updated progressively for scalability. We want to guarantee interactive latency while avoiding PVA from interfering with the user’s exploration process.
3.2.2 Control Panel
{subcaptiongroup}
Figure 3:
Figure 4:
Figure 5: 3(a): The “Detail” tab. 5: The “Filtering” tab. 5: The “History” tab. See Visualization of Electronic Health Record Sequences at Scale for the Overview tab.
Overview The “Overview” tab (Visualization of Electronic Health Record Sequences at Scale) shows (1) the number of high-level events aggregated and (2) the hierarchy associated with the data using an Icicle tree. Hovering a node shows the number of aggregated events it contains. Those numbers are updated continuously when the progressive algorithm is running.
The hierarchy groups multiple high-level events into a higher category, e.g., all drugs under the category “medication”. Visualization of Electronic Health Record Sequences at Scale visualizes the hierarchy and its distribution in the partial aggregated tree. For applications with many event types, using higher hierarchy levels simplifies the visualization. Users can modify the level of hierarchy at any time. It is instantaneous, as it only impacts the prefix tree visualization without altering how the progressive algorithm aggregates the low-level events.
Detail The “Detail” tab (3(a)) shows the data distributions of a selected node. Currently, we show the age and duration distributions using histograms and the comorbidities using a bar chart. Those graphs are updated iteratively when the progressive algorithm runs.
Filtering The “Filtering” tab (Figure 5) allows to select a subset of patients and to filter the view according to parameters and node attributes. When users adjust the aggregating parameters (e.g., the thresholds defining the interruption events) or select a subset of patients by specifying query parameters, the progressive algorithm restarts from the beginning, considering only sequences matching the query and updating the rules that aggregate low-level events into high-level events. In contrast, filtering the view only manipulates nodes from the already computed tree (e.g., hide small nodes or merge nodes into higher hierarchical levels); it does not process the data again and results in instantaneous updates. To filter the view, users can specify a hierarchical level, truncate the tree to a maximum depth (e.g., show only the first three depth levels), and select a minimum number of patients per sequence to hide small nodes, simplifying the tree.
History and Comparison The “History” tab (Figure 5) is inspired by the history comparison feature of Pister et al. [36] that relies on the Trrack [10] history management library. It shows the history of the user’s actions as a tree; the user can jump to any past state of the exploration by selecting a history node.
The user can compare the current state with another one using the “compare to” button. ParcoursVis initializes a second context for the second state. The user can visualize the aggregated trees from both contexts in a back-and-forth manner using the “Next” and “Previous” buttons (not shown on the screenshots). The node associated with the current state in the “History” tab has a blue halo, while the compared node has a green halo. Selecting this second context initially restarts the progressive aggregation using the parameters associated with its state. As the context being compared may not have finished processing all patients when switching back and forth between the compared contexts, we do not restart the aggregation from the beginning but resume it from where it was left previously. This avoids unnecessary waiting for the user to get back to where the context was previously and avoids breaking the user’s mental model related to the context because the progressive algorithm may never get back to the exact previous view.
Overall, the PVA aspect of ParcoursVis required us to consider the usability issues of every feature that either (1) forces the aggregation algorithm to restart, (2) conversely, avoids restarting the algorithm when the front end can handle it alone, and (3) affects the user experience in an unexpected way (e.g., as with the ”History“ tab).
3.3 Implementation
CSV file
[…]
Figure 6: Architecture of ParcoursVis, organized in a pre-processing stage, a back end, and a front end. See the state of the tree in red.
Figure 6 describes the architecture of ParcoursVis. Our front-end application is made of about 3000 lines of JavaScript that uses D3.js [7] and Vue.js [47]. This allows us to show visualizations on a regular web browser accessible to anyone and, in particular, to referring doctors without any installation. However, web technologies are not designed for heavy computation. We, therefore, rely on an optimized back-end library written in C++ (3000 lines) and used as a Python module that progressively aggregates the data sent to the front end. The back-end Python server code is 127 lines long. We also provide test suites for JavaScript and C++ that double the running code size. The code is open-source at https://gitlab.inria.fr/aviz/parcoursvis/.
3.4 Optimization of ESS Visualization Programs
@@
First, as shown in Figure 6, ParcoursVis relies on pre-parsed data stored in binary format, avoiding the expensive step of parsing CSV data at each program start while ensuring all events per patient are sorted chronologically and that patients are shuffled to optimize the aggregation.
Running the preprocessing step takes three minutes for 50 million events stored in CSV format (as in Table 1). It is performed once for a given dataset and generates the binary files that are memory-mapped when ParcoursVis’s back end starts.
Whereas the non-progressive program loads the CSV files in a time linear to the file size (LABEL:lst:simple line 2), the progressive back end loads the binary files in negligible time using memory mapping, as shown in LABEL:lst:aggregation lines 3–4.

### v2

# ParcoursVis: Visualization of Electronic Health Record Sequences at Scale

@@
Abstract.
We present ParcoursVis, an open-source Progressive Visual Analytics tool designed to explore aggregated electronic health record sequences of patients at scale. Existing tools are limited to about 20k patients that they can process fast enough to remain interactive, under human latency limits. They need to process the whole dataset before showing the visualization, taking a time proportional to the data size. Yet, managing large datasets allows for discovering rare medical conditions and unexpected patient pathways, contributing to improving treatments. To overcome this limitation, ParcoursVis relies on a progressive aggregation algorithm that quickly computes an approximate initial result, visualized as an Icicle tree, and improves it iteratively, until the whole computation is done. With its architecture, ParcoursVis remains interactive while visualizing the sequences of millions of patients—three orders of magnitude more than similar tools. We describe our PVA architecture, which achieves scalability with fast convergence and visual stability.
@@ Abstract.
copyright: rightsretainedjournalyear: 2026conference: CHI Conference on Human Factors in Computing Systems; April 13-17, 2026; Barcelona, Spainbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13-17, 2026, Barcelona, Spaindoi: 10.1145/XXXXXXX.XXXXXXXisbn: XXX-X-XXXX-XXXX-X/XX/XXccs: Human-centered computing Visualization systems and toolsccs: Information systems Data management systems
@@
Figure 1. The ParcoursVis system. (1) The main view shows 10 M medical pathways aggregated, visualized as an Icicle tree, for patients treated for non-cancerous prostate adenoma. (1) The number of patients visible and filtered out. (1) The control panel with the legend and overall distribution of event types (e.g., buying medication or having surgery). (1) The hierarchy of event types, used to select a level of detail of the tree. (1) Zoomed view of a sub-tree, revealing more of the hierarchical structure. Each colored node represents a stage in a sequence of events starting from the root (left), flowing to the right. Child nodes representing the next stages are on the right of their parent, sorted by frequency (i.e., proportion of patients) from bottom to top. Node height is proportional to the frequency of the event, and width to the average duration (in days). Here, most patients started with the alphabloc treatment (bottom left). :
@@
[…]
1. Introduction
Visualization scalability is particularly challenging in healthcare, where Electronic Health Records (EHRs) can contain a large amount of multivariate temporal data for millions of patients over decades. As datasets increase in size and complexity, traditional visualization techniques struggle to maintain interactivity and responsiveness, thereby hindering effective exploratory data analysis on these large datasets. Indeed, results from research in visualization show that user attention declines when latency exceeds 500 ms, and after five to ten seconds, users tend to abandon tasks or lose focus (Zgraggen et al., 2017; Liu and Heer, 2014). Addressing the scalability challenge necessitates improvements in algorithms and the development of optimized data structures within visualization systems, but even with the heaviest optimizations (Heer and Moritz, 2024; Liu et al., 2013), existing visualization systems’ latency remains proportional to dataset size, limiting their scalability. To overcome this limitation, Progressive Visual Analytics (PVA) appears as a novel programming paradigm that decouples data size from latency (Fekete et al., 2024) by quickly computing and visualizing an approximate initial result and improving it iteratively until the whole computation is done. Users can interact with the visualization at any step with low latency. PVA has been effectively applied to several application domains (Ulmer et al., 2024), but it introduces new challenges, such as a new programming architecture, as well as visualization stability, uncertainty management, and potentially usability issues, that should be considered for each new application, and in particular for the exploration of large EHR datasets.
For instance, the SNDS (the French national health database (SNDS, 2023)) holds decades of health data on all the French citizens, encompassing records of drug purchases, medical diagnoses, and hospital treatments. Exploring such data has the potential to advance healthcare by enabling a better understanding of treatment practices and patient outcomes. This approach aligns with the goal of medical professionals to assess whether patients’ treatments adhere to national guidelines and, if deviations occur, to understand the reasons. This is key to providing patients with the highest quality of care through evidence-based medicine and verified recommendations. Yet, the volume of such datasets necessitates innovative visualization approaches capable of handling large medical data while preserving efficiency and usability.
In this context, we developed ParcoursVis (Figure 1), a PVA tool created in collaboration with medical institutions to visualize patient care pathways derived from EHRs. Previous work has laid a foundation for similar visualization tools aimed at healthcare analysis (e.g., (Bjarnadóttir et al., 2016; Meyer et al., 2013; Ozkaynak et al., 2015)). For instance, EventFlow (Monroe et al., 2013) aggregates event sequences into a prefix tree and visualizes it as an Icicle tree, as shown in Figure 1. This visualization has proven effective for several applications (e.g., (Bjarnadóttir et al., 2016; Monroe et al., 2013; Ozkaynak et al., 2015)). We, therefore, build on the EventFlow visualization technique for ParcoursVis, along with several improvements.
However, tools like EventFlow have generally been applied to relatively small datasets, sometimes encompassing fewer than 100 patients. According to our experiments, EventFlow remains interactive up to about 20 k patients. While such tools have offered valuable initial insights, they fall short in scalability, which has been reported as a main challenge for EHR visualization (Wang and Laramee, 2022; Arleo et al., 2024). In practice, regional and national EHR databases present data at a far greater scale, often involving millions of patients. To our knowledge, no existing tool enables the visualization and interactive analysis of more than about 100 k EHR aggregated event sequences (Wongsuphasawat et al., 2011; Wongsuphasawat and Gotz, 2012; Monroe et al., 2013; Wang and Laramee, 2022), which limits the scope of exploration. This constraint impedes the ability to gain a more comprehensive view of patient pathways, identify patterns and trends that may not be visible in smaller or sampled datasets, and discover rare patterns and outliers.
With ParcoursVis, we allow visualizing temporal event sequences of more than 100 M patients, potentially supporting the EHRs of the largest countries. Our progressive approach provides a quick, approximate, yet accurate preview of the data early on and quick updates thereafter, converging within seconds. After the initial preview, our algorithm continues processing the dataset, updating the visualization every second or so until it is fully processed.
Yet, the progressive updates of the aggregated tree can cause visual instability, as the nodes (i.e., rectangles on Figure 1) are laid out at each update. The shifting of nodes can disrupt the viewer’s mental map and create distracting visual flickers. Therefore, our approach may have an impact on the usability of the tool. To mitigate this issue, we introduce a sorting with hysteresis algorithm that limits this flicker during progressive rendering.
In this work, we describe our PVA architecture and explain the changes to turn a traditional architecture for visualizing aggregated sequence data into a progressive one.
In summary, this article presents our system and reports on its efficiency in terms of PVA, stability, and possibly flicker. Specifically:
  1. (1)
    We describe our PVA architecture to achieve scalability for aggregated event-sequence visualization over the web.
  2. (2)
    We introduce our algorithm, specifically designed to process events progressively in order to provide meaningful intermediate results, and we evaluate its efficiency using metrics specific to PVA (Richer et al., 2024b). In particular, we report on scalability, and latency.
  3. (3)
    To optimize visual stability despite the potential flicker introduced by PVA, we introduce a sorting algorithm with hysteresis, i.e., maintaining node order when size differences remain below a threshold. We evaluate stability through quantitative measures, showing that hysteresis-based sorting improves both stability and convergence of the progressive visualization.
  4. (4)
    We report on usability regarding the progressive aspect. Our professional participants have completed a variety of tasks seamlessly, without mentioning any delays in the system’s response, or discomfort with the progressive rendering of the Icicle tree.
We address a key challenge in aggregated EHR sequence visualization and present ParcoursVis wich improves scalability by at least three orders of magnitude compared to state-of-the-art systems. Central to this contribution, we present the first PVA tree visualization algorithm (according to the survey of Ulmer et al. (Ulmer et al., 2024)), and building on the results of PVA systems in other domains (Ulmer et al., 2024), we show that ParcoursVis maintains interactivity, usability, and utility while ensuring scaling to much larger datasets. Ultimately, our goal is for ParcoursVis to provide a user experience comparable to EventFlow, but capable of handling up to 10,000 times more patients without users perceiving the progressive nature of the system.The source code and evaluation scripts are available at https://gitlab.inria.fr/aviz/parcoursvis/.
2. Related Work
ParcoursVis reuses the visualization technique of EventFlow because it was thoroughly tested and validated for analyzing medical temporal event sequences (e.g.,  (Bjarnadóttir et al., 2016; Monroe et al., 2013; Ozkaynak et al., 2015)). Thus, our research goal is not to propose a new visualization technique but rather to explore the strategies to scale and adapt existing approaches that have already demonstrated their effectiveness and apply them to the largest EHR data available. Especially, we consider the usability of EventFlow’s visualization validated and only evaluate ParcoursVis for the specific features it introduces, specifically the use of a progressive algorithm to ensure scalability and our strategy to avoid subsequent instability. Therefore, this work relates to temporal event sequences, scalability, PVA systems, and progressive uncertainty in the partial results.
2.1. Temporal Event Sequences